OpenStudy (anonymous):

tan^4theta-sec^4theta smiplify

OpenStudy (gorv):

1+2sec^2theta 0r 2tan^2theta-1

OpenStudy (anonymous):

\[\tan ^{4}\theta-\sec ^{4}\]

OpenStudy (gorv):

(tan^2-sec^2)*(tan^2+sec^2) tan^2-sec^2=1 so ans = tan^2+sec^2

OpenStudy (gorv):

@Dixiegirl4eve got it????

OpenStudy (anonymous):

my answer is way off